Skyler is 20 years older than Jason. In 8 years, Skyler will be twice as old as Jason. How old is each now?
Answer by Maths68(1474)

Let
Present age of Skyler = s
Present age of Jason = j
Given
Skyler is 20 years older than Jason
s=j+20.................(1)
In 8 years
Skyler will be = s+8
Jason will be = j+8
Then
Skyler will be twice as old as Jason
s+8=2(j+8)
s+8=2j+16
s=2j+16-8
s=2j+8..............(2)
Put the value of s from (1) to (2)
j+20=2j+8
j-2j=8-20
-j=-12
Multiply by -1 both sides
(-1)(-j)=(-1)(-12)
j=12
Put the value of j in (1)
s=j+20
s=12+20
s=32

Present age of Skyler = s = 32 years old
Present age of Jason = j = 12 years old